OpenStudy (anonymous):

The area of a rectangle is 4,320 ft2. The ratio of the length to the width is 6:5. Find the length of the rectangle

OpenStudy (anonymous):

Let x = length and y = width The ratio of the length to the width is 6:5 This tells us that x / y = 6 / 5 The area of a rectangle is 4,320 What formula can you write given that sentence..involving x and y?

OpenStudy (campbell_st):

why not just make it simple... let x be a measurement then the dimensions of the rectangle are length = 6x and width = 5x then \[6x \times 5x = 4320\] just solve for x. hope it helps.